OpenStudy (anonymous):

Mark has $4.50 and Jean-Martin has $3.00. Mark spends twice as much as Jean-Martin and now sees that he has half as much money left as Jean-Martin has left. How much money did Mark spend? How much money did Jean-Martin spend? How much does each of them have left?

OpenStudy (mertsj):

Let x = amount Jean-Martin spends. 4.5-2x = amount Mark has left 3-x=amount Jean-Martin has left. Mark has 1/2 the amount of Jean-Martin 4.5-2x=1/2(3-x)